# Efficient Stochastic Template Bank using Inner Product Inequalities

### v1

@@ 1 Introduction
Various techniques exist for constructing a template bank. The two main classes of methods are geometric (Brown et al., 2012; Babak et al., 2006; Schmidt et al., 2024), stochastic placement techniques (Harry et al., 2009; Ajith et al., 2014; Manca & Vallisneri, 2010; Babak et al., 2006). There also exist hybrid methods that utilizes both approaches. (Roy et al., 2017, 2017). Geometric methods typically place template waveforms using a lattice technique; this generally requires a known metric, which directly quantifies the similarity of signals. Afterwards, the templates are converted into a physical parameter space. This approach guarantees that every template captures a predefined minimum percentage of the signal (Owen, 1996; Prix, 2007). One challenge is that the metric may only be flat in the non-physical space. This may result in over-placing templates near and outside the physical space boundaries. These templates need to be included since they may be better at recovering signals near the boundary than the templates within the physical space (Cokelaer, 2007; Coogan et al., 2022; Brown et al., 2012). Consequently, since a search’s computational cost is proportional to the number of templates produced from any method, including these templates can lead to an excess in the template count, making the search more expensive to perform. This approach is not optimal for complicated boundaries, unknown metrics, and for more intricate detection studies requiring non-trivial parameter spaces.
@@
In this paper, we present a new stochastic placement method. We show this method produces self-consistent banks that recover the SNR of any potential signal within a target search space, with losses below a chosen threshold. We show that this method has already been used to conduct various searches such as (low-mass and eccentric) neutron stars, (supermassive and primordial) black holes, and neutron-star–black-hole binaries that produce gamma-ray bursts. Finally, we show how the number of parameters in a bank scales the time to complete a search by generating different banks with varying intrinsic parameters such as spin, eccentricity, and or tidal deformability.
@@ 2 Evaluation of Template Bank Coverage
Where $S_{n}(f)$ is the power spectral density and $f$ is frequency. Since an overall phase and the absolute time of arrival are typically nuisance parameters in a search, we define the phase and time maximized overlap to be the match $\mathcal{M}$. This formalism is for a non-precessing signal. For precessing signals, refer to Harry et al. (2016).
@@ 3 Methods: Stochastic Template Bank Algorithm
The bank is sufficiently covered under two termination criteria: minimal match and tolerance. Tolerance is the fraction of accepted templates over the total amount of templates in the bank. This criterion determines how many templates must be drawn to sufficiently cover the bank. Lower tolerance values generate banks with more templates to cover the parameter space adequately; larger tolerance values do the opposite. The minimal match is the minimum percentage of SNR that at least one template in the bank can recover any fiducial signal. We use this criterion to accept or reject proposal templates. If the match between the proposal and the accepted templates is less than the minimal match, that proposed template will be added to the bank. If the template has a match greater than the minimal match condition, sufficient templates exist in the bank to recover any given signal in that region.
@@ 3.1 Template Sampling Algorithm
To define the $\tau_{0}$ boundaries, we calculate the chirp time defined above using a default lower chirp time frequency of 15Hz for the minimum and maximum mass range corresponding to the start and end $\tau_{0}$. The start $\tau_{0}$ will correspond to the highest masses in the space and the end will correspond to the lowest mass pair. The crawl parameter determines the width of each bin. Once the $\tau_{0}$ boundaries are defined, we draw the first set of points from the prior distribution of the parameters within the first $\tau_{0}$ bin. The sampled templates are either accepted or rejected into the bank (elaborated in section 3.2). Sampling all the templates randomly can result in banks with a lot of holes, or lacking templates. To combat some of the gaps in the bank, we utilize a Gaussian Kernel Density Estimate (KDE). Once sufficient templates accumulate in the bank, we generate a probability density function (PDF) from the accepted proposals. From the PDF, we can identify poor regions where there are divets or tails corresponding to lack of templates and propose templates to aid in sufficiently covering the sparser regions. Another benefit for utilizing two different sampling methods is the ability to speed up proposals when one method performs worse than the other. When the fraction of accepted templates over all proposed templates, the tolerance, is worse for either sampling method, the technique will swap to drawing proposals using the alternate technique. Once the first $\tau_{0}$ bin has been sufficiently filled with templates and the tolerance criteria is achieved, we begin sampling in the next $\tau_{0}$ bin overlapping 50% with the previous strip. Figure 2 shows how the $\tau_{0}$ strips overlap with each bin as the sampler proceeds to cover the entire parameter space.
@@ 3.2 Optimizing Template Proposal Acceptance
Once a template is proposed, we must determine if it should be added to the template bank to enhance its coverage of the target parameter space. This is done by determining the FF of our proposed template against the current set of templates in the bank to assess the similarity. Templates with FFs less than the minimal match condition are not sufficiently similar and need to be added to the bank. Templates exceeding the minimal match condition are similar and do not need to be included in the bank. To start, we generate a reduced frequency waveform model, inspired by the methodology detailed in Ajith et al. (2014); Capano et al. (2016). This model significantly reduces the computational load required for generating waveforms for the match calculations. Next, we evaluate the matches between the proposed waveform models and the closest templates in chirp time, storing these values. To rule out templates where the match is clearly above the minimal match criterion, we make use of previously stored matches. We further optimize this procedure using triangle inequality which states that the sum of any two sides of any triangle will be greater than the third side. If we compare the distances between templates, we will automatically know that the third distance between the templates will be smaller. For example, consider the mismatch (1 - $\mathcal{M}$), which represents the distance of the templates, between waveform A and waveform B, and between waveform B and waveform C. If the sum of mismatches between A and B, B and C is smaller than the predefined 1-FF threshold, we can skip the direct calculation of the match between A and C.
@@ 4 Bank Verification
Table 1: The computational time to generate template banks and conducting a search for different parameter spaces. The bolded parameters corresponds to template banks generated using our method. The first three banks share the same mass-only parameter space but differ in the methods used: the geometric bank (Brown et al., 2012), an existing stochastic method (Ajith et al., 2014; Capano et al., 2016), and our method respectively. The last three rows use our method to construct template banks with spin, eccentricity and tidal deformability parameters. For searches, we assume one core processes 5000 templates in real time and estimate the time required to analyze one year of data from three detectors. The time it takes to complete a search is a factor of $\mathbf{O}(10^{5})$ more than the time to complete the bank generation for the non-spinning and only spin aligned case, $\mathbf{O}(10^{4})$ for the eccentric bank, and $\mathbf{O}(10^{3})$ for the tidal deformable case. The non-spinning geometric technique generates banks faster than the stochastic methods for mass only banks. However, the time ratio is the largest for geometric method implicating that currently template bank generations are a negligible cost to running a search, and faster search methods should be investigated. For larger parameter spaces, the banks generated in around 30 to 60 core days. The fraction of search time over bank generation time was the lowest for these scenarios. Overall, the most time consuming process is the search. The cost of generating the banks are trivial to the cost of running the search and faster search methodologies should be developed.
@@
$\boldsymbol{m_{1}},\boldsymbol{m_{2}},\boldsymbol{\chi_{1}},\boldsymbol{\chi_{2}},\boldsymbol{\lambda_{1}},\boldsymbol{\lambda_{2}}$ $1.86\times 10^{6}$ 55.697 $4.08\times 10^{5}$ $7.32\times 10^{3}$ TaylorF2
5 Computational Scaling
@@
As we increase the number of parameters, we see the banks take a longer time to generate and the number of templates increase by roughly $\mathcal{O}(10)$ for every additional pair of parameter incorporated. Explicitly looking at the TaylorF2 waveform banks, mass only had $\mathcal{O}(10^{4})$ and adding a pair of spin parameters drove the bank size up to $\mathcal{O}(10^{5})$ and the tidal deformability banks was of size $\mathcal{O}(10^{6})$. The eccentric bank and the tidal deformability bank took the longest to generate. The eccentric bank was faster by 23.46 core days, but had 3.04 times more templates. One reason for this difference in bank size could be due to the choice of priors. If we increase the priors to a larger range, there are more options of templates proposals and more templates are necessary to sufficiently cover the parameter space. However, the excess number of templates does not seem to limit the computational speed in the eccentric bank. One explanation for this difference can be due to the use of different waveform approximants. The eccentric waveform might have been faster in producing waveforms such that the template comparisons are not limited by the approximant generation cost.
@@ 6.5 Open Gravitational-Wave Catalog Search for Compact-Binary Mergers
The fourth open gravitational-wave catalog (OGC) contains the observation of nearly 100 compact binary mergers (Nitz et al., 2023). The search covers a parameter space from neutron star binaries through heavy binary black holes (up to $\sim 1000$ $M_{\odot}$). The OGC search search splits the analysis into four sub-regions, covering BNS, NSBH, BBH, and a focused BBH region where the largest numbers of signals are observed. The analysis based on the open source PyCBC toolkit (Usman et al., 2016; Davies et al., 2020) and makes use of the flexible template bank generation introduced in this work. This enables the use of multiple waveform approximants depending on the suitability for different parts of parameter space, e.g. the use of TaylorF2 (Sathyaprakash & Dhurandhar, 1991b; Droz et al., 1999; Blanchet, 2002; Faye et al., 2012) for neutron star binaries, and IMRPhenomD (Husa et al., 2016; Khan et al., 2016b) and SEOBNRv4 for BBH signals (Taracchini et al., 2014; Bohé et al., 2016). The stochastic placement algorithm allows for the iterative generation of a larger template bank by adding to a pre-existing template bank. This allows for regions of parameter space that adapt to different requirements, e.g. a higher minimal match for the focused BBH region of 0.995 or the inclusion of tidal deformability for binary neutron star templates.

### v2

@@ 1 Introduction
Various techniques exist for constructing a template bank. The two main classes of methods are geometric (Brown et al., 2012; Babak et al., 2006; Schmidt et al., 2024), stochastic placement techniques (Harry et al., 2009; Ajith et al., 2014; Manca & Vallisneri, 2010; Babak, 2008). There also exist hybrid methods that utilizes both approaches. (Roy et al., 2017, 2017). Geometric methods typically place template waveforms using a lattice technique; this generally requires a known metric, which directly quantifies the similarity of signals. Afterwards, the templates are converted into a physical parameter space. This approach guarantees that every template captures a predefined minimum percentage of the signal (Owen, 1996; Prix, 2007). One challenge is that the metric may only be flat in the non-physical space. This may result in over-placing templates near and outside the physical space boundaries. These templates need to be included since they may be better at recovering signals near the boundary than the templates within the physical space (Cokelaer, 2007; Coogan et al., 2022; Brown et al., 2012). Consequently, since a search’s computational cost is proportional to the number of templates produced from any method, including these templates can lead to an excess in the template count, making the search more expensive to perform. This approach is not optimal for complicated boundaries, unknown metrics, and for more intricate detection studies requiring non-trivial parameter spaces.
@@
In a broader scientific context, the construction of template banks can be likened to the principles of experimental design, where the goal is to efficiently sample a high-dimensional parameter space, similar to methods used in statistics, computer science, and manufacturing for uncertainty quantification and optimization (Sacks et al., 1989). Each experiment in these fields corresponds to a specific set of input parameters, and the objective is often to understand the variation of a Quantity of Interest (QoI) based on these parameters, ensuring adequate space-filling to capture the underlying behavior of the system (Sacks et al., 1989). Similarly, template banks for gravitational wave detection aim to cover the parameter space as fully as possible to maximize detection efficiency and minimize missed signals.
In this paper, we present a new stochastic placement method. We show this method produces self-consistent banks that recover the SNR of any potential signal within a target seqarch space, with losses below a chosen threshold. We show that this method has already been used to conduct various searches such as (low-mass and eccentric) neutron stars, (supermassive and primordial) black holes, and neutron-star–black-hole binaries that produce gamma-ray bursts. Finally, we show how the number of parameters in a bank scales the time to complete a search by generating different banks with varying intrinsic parameters such as spin, eccentricity, and or tidal deformability.
@@ 2 Evaluation of Template Bank Coverage
Where $S_{n}(f)$ is noise power spectral density and $f$ is frequency. Since an overall phase and the absolute time of arrival are typically nuisance parameters in a search, we define the phase and time maximized overlap to be the match $\mathcal{M}$. This formalism is for a non-precessing signal. For precessing signals, refer to Harry et al. (2016).
@@ 3 Methods: Stochastic Template Bank Algorithm
The bank is sufficiently covered under two termination criteria: minimal match and tolerance. Tolerance is the termination criteria threshold defined by the user. When the tolerance falls below the acceptance fraction, the fraction of accepted versus the proposed templates, the region of a template bank is considered to be completed. This criterion determines how many templates must be drawn to sufficiently cover the bank. Lower tolerance values generate banks with more templates to cover the parameter space adequately; larger tolerance values do the opposite. The minimal match is the minimum percentage of SNR that at least one template in the bank can recover any fiducial signal. We use this criterion to accept or reject proposal templates. If the match between the proposal and the accepted templates is less than the minimal match, that proposed template will be added to the bank. If the template has a match greater than the minimal match condition, sufficient templates exist in the bank to recover any given signal in that region.
@@ 3.1 Template Sampling Algorithm
To define the $\tau_{0}$ boundaries, we calculate the chirp time defined above using a default lower chirp time frequency of 15Hz for the minimum and maximum mass range corresponding to the start and end $\tau_{0}$. The start $\tau_{0}$ will correspond to the highest masses in the space and the end will correspond to the lowest mass pair. The crawl parameter determines the width of each bin. Once the $\tau_{0}$ boundaries are defined, we draw the first set of points from the prior distribution of the parameters within the first $\tau_{0}$ bin. The sampled templates are either accepted or rejected into the bank (elaborated in section 3.2).
The objective of an effective stochastic proposal is to ensure comprehensive coverage of the parameter space, minimizing gaps and undercoverage. An ideal stochastic proposal would be one proportional to the final template density, which is uniform in metric space. In general, one may not know this distribution a priori. To approximate the ideal distribution, and minimize the chance of gaps, holes, or undercoverage, we utilize a Gaussian Kernel Density Estimate (KDE). We construct the KDE from the most recently accepted template parameters after a sufficiently large population is collected using a chosen prior proposal distribution. The KDE distribution updates as new points are accepted and can then evolve towards the ideal distribution. However, the KDE alone cannot fully resolve issues arising from complex boundaries within the parameter space. Hence, the algorithm switches between drawing uniformly from the prior proposal and the KDE-based proposal based on which is more efficient at the time. Once the first $\tau_{0}$ bin has been sufficiently filled with templates and the tolerance criteria is achieved, we begin sampling in the next $\tau_{0}$ bin overlapping 50% with the previous strip. Figure 2 shows how the $\tau_{0}$ strips overlap with each bin as the sampler proceeds to cover the entire parameter space.
To parallelize the bank generation, we divide the parameter space into sub-banks, each of which can be processed independently. To achieve efficient parallelization, each sub-bank is run over a set of independent $\tau_{0}$ bins, chosen so that the match between templates across boundaries of adjacent sub-banks remains below the minimal match threshold. Because templates are compared across $\tau_{0}$ bins within the generation of each sub-bank, over-coverage only occurs in the bounding $\tau_{0}$ bins as they are not compared to templates across shared boundaries with other sub-banks. This means that if each sub-bank has 50 $\tau_{0}$ bins, then the over-coverage is limited to less than $\sim 2\%$. For most analyses, we consider this excess to be negligible.
@@ 3.2 Optimizing Template Proposal Acceptance
Once a template is proposed, we must determine if it should be added to the template bank to enhance its coverage of the target parameter space. This is done by determining the FF of our proposed template against the current set of templates in the bank to assess the similarity. Templates with FFs less than the minimal match condition are not sufficiently similar and need to be added to the bank. Templates exceeding the minimal match condition are similar and do not need to be included in the bank. To start, we generate a reduced frequency waveform model, inspired by the methodology detailed in Ajith et al. (2014); Capano et al. (2016). This model significantly reduces the computational load required for generating waveforms for the match calculations. PSD fluctuations that typically occur during an observing run have been found to have minimal impact on template placement and coverage; standard searches use a template bank generated on the expected or average noise curve of an observing run  (Usman et al., 2016). Next, we evaluate the matches between the proposed waveform models and the closest templates in chirp time, storing these values. To rule out templates where the match is clearly above the minimal match criterion, we make use of previously stored matches. We further optimize this procedure using triangle inequality which states that the sum of any two sides of any triangle will be greater than the third side. If we compare the distances between templates, we will automatically know that the third distance between the templates will be smaller. For example, consider the mismatch (1 - $\mathcal{M}$), which represents the distance of the templates, between waveform A and waveform B, and between waveform B and waveform C. If the sum of mismatches between A and B, B and C is smaller than the predefined 1-FF threshold, we can skip the direct calculation of the match between A and C.
@@ 4 Bank Verification
5 Computational Scaling
Table 1: The computational time to generate template banks and conduct a search for different parameter spaces. The bolded parameters correspond to template banks generated using our method. The first three banks share the same mass-only parameter space but differ in the methods used: the geometric bank (Brown et al., 2012), an existing stochastic method (Ajith et al., 2014; Capano et al., 2016), and our method respectively. Our method produces comparable sized banks to the other methods in similar times. However, unlike the geometric and existing stochastic method, our method can generate banks with higher dimensional spaces that include eccentricity, deformability and so on. The last three rows use our method to construct template banks with spin, eccentricity and tidal deformability parameters. For searches, we assume one core processes 5000 templates in real time and estimate the time required to analyze one year of data from three detectors. The time it takes to complete a search is a factor of $\mathbf{O}(10^{5})$ more than the time to complete the bank generation for the non-spinning and only spin aligned case, $\mathbf{O}(10^{4})$ for the eccentric bank, and $\mathbf{O}(10^{3})$ for the tidal deformable case. The non-spinning geometric technique generates banks faster than the stochastic methods for mass only banks. However, the time ratio is the largest for geometric methods, implying that currently template bank generations are a negligible cost to running a search, and faster search methods should be investigated. For larger parameter spaces, the banks generated in around 30 to 60 core days. The fraction of search time over bank generation time was the lowest for these scenarios. Overall, the most time consuming process is the search. The cost of generating the banks is trivial to the cost of running the search and faster search methodologies should be developed.
@@ 5 Computational Scaling
$\boldsymbol{m_{1}},\boldsymbol{m_{2}},\boldsymbol{\chi_{1}},\boldsymbol{\chi_{2}},\boldsymbol{\lambda_{1}},\boldsymbol{\lambda_{2}}$ $3.94\times 10^{6}$ 29.21 $8.63\times 10^{5}$ $2.96\times 10^{4}$ TaylorF2
@@
As we increase the number of parameters, we see the banks take a longer time to generate and the number of templates increase by roughly $\mathcal{O}(10)$ for every additional pair of parameter incorporated. Explicitly looking at the TaylorF2 waveform banks, mass only had $\mathcal{O}(10^{4})$ and adding a pair of spin parameters drove the bank size up to $\mathcal{O}(10^{5})$ and the tidal deformability banks was of size $\mathcal{O}(10^{6})$. The eccentric bank and the tidal deformability bank took the longest to generate and had the most amount of templates.
We find that our method performs similarly to existing stochastic and geometric methods in generating a two-parameter mass-only bank. The geometric method produced the bank in the shortest time and the second largest bank. The existing stochastic method produced the most templates, but also took the longest to generate a bank. However, the comparable cost to complete the search was the lowest for sbank. Overall, our method performs just as well as existing methods. The main advantage of our method is its ability to efficiently construct high-dimensional parameter spaces ($N>4$), in which these other methods are unable to do so. Utilizing the inner product inequality significantly enhances efficiency, particularly for larger, higher-dimensional parameter spaces. To quantify the impact of this efficiency, we estimate the proportion of matches excluded by the inner product inequality. For instance, in one of the largest tidal deformable sub-banks, we observed a 40$\sim$ 40$\times$ 40 × reduction in the number of required matches.
@@ 6.5 Open Gravitational-Wave Catalog Search for Compact-Binary Mergers
The fourth open gravitational-wave catalog (OGC) contains the observation of nearly 100 compact binary mergers (Nitz et al., 2023). The search covers a parameter space from neutron star binaries through heavy binary black holes (up to $\sim 1000$ $M_{\odot}$). The OGC search search splits the analysis into four sub-regions, covering BNS, NSBH, BBH, and a focused BBH region where the largest numbers of signals are observed. The analysis based on the open source PyCBC toolkit (Usman et al., 2016; Davies et al., 2020) and makes use of the flexible template bank generation introduced in this work. This enables the use of multiple waveform approximants depending on the suitability for different parts of parameter space, e.g. the use of TaylorF2 (Sathyaprakash & Dhurandhar, 1991b; Droz et al., 1999; Blanchet, 2002; Faye et al., 2012) for neutron star binaries, and IMRPhenomD (Husa et al., 2016; Khan et al., 2016b) and SEOBNRv4 for BBH signals (Taracchini et al., 2014; Bohé et al., 2017). The stochastic placement algorithm allows for the iterative generation of a larger template bank by adding to a pre-existing template bank. This allows for regions of parameter space that adapt to different requirements, e.g. a higher minimal match for the focused BBH region of 0.995 or the inclusion of tidal deformability for binary neutron star templates.